OpenStudy (anonymous):

Can someone help me with physics. I tried, but I have no idea what to do. My values come out wrong. 1)A bucket is filled to the brim with water. A rock is dropped into the bucket, causing 75.7 mL of water to spill out. Once dropped into the bucket, the rock appears to have 82.0 % of its original mass. Use this information, along with Archimedes principle, to determine the true mass of the rock in grams. Note: The density of water is 1.00 g/cm3, and 1 L = 1000 cm3.

OpenStudy (anonymous):

2)Water flows through a rectangular pipe at a speed of 1.24 m/s. The cross-sectional area of the pipe has a width W = 1.61 cm and a length L = 26.0 cm. The pipe then runs through a connection that changes the cross sectional area of the pipe to a circle. If the speed of water in the circular pipe is now 23.0 m/s, what is the radius R of the new pipe? Give your answer in cm. Note: ignore viscosity.

OpenStudy (anonymous):

3)A garden hose is used to fill a large metal container which can hold 21.0 L of liquid. If the radius of the garden hose's nozzle is 1.1 cm, and the speed of the water at this point is 273.0 cm/s, how long (in seconds) would it take to fill the container with water, assuming that none of it splashes out? Hint: Use the fact that 1 L = 1000 cm3

OpenStudy (anonymous):

4)A block of birch floats in oil with 94.0% of its volume submerged. What is the density of the oil (in g/cm3)? The density of the birch is 0.56 g/cm3.

OpenStudy (anonymous):

anyone of the questions is fine. just help

OpenStudy (anonymous):

If the density of the oil equaled that of the birch, then the entire volume of the birch would be submerged (barely float). If the oil had half the birch density, the block would have half its volume submerge. Here 94% is submerged so the oil has a density of 0.94(0.56g/cm3)= 0.526 g/cm3.

OpenStudy (anonymous):

(3) The rate of volume flow, Q in cm3/s equals the average velocity (cm/s) times the cross-sectional area, (pi/4)R^2. Q = v A. The time it takes to fill the volume V is just t=V/Q. Take it from there.

OpenStudy (anonymous):

(2) The volume rate of flow is Q = v A, velocity times cross-sectional area. In the first case, A is rectangular length x width, and in the second it is the area of the circular opening (pi/4) R^2.

OpenStudy (anonymous):

(1) The amount of water spilled is the volume of the rock. The weight of the water decreases the apparent weight of the rock by just that amount. Must go.

OpenStudy (anonymous):

Oops. The cross-section area is pi R^2 not pi/4 R^2. Sorry! Haste makes waste.

OpenStudy (anonymous):

thank you. could u explain 1

OpenStudy (anonymous):

i dont get what u r saying

OpenStudy (anonymous):

Which question is puzzling? Volume rate of flow, Q equals the average velocity times the cross-sectional area of the opening, length times width for the rectangle pi times the radius squared for the circular opening. The time it takes to fill the volume is t = V/Q.